OpenStudy (usukidoll):

Let \(X\) be a set and consider the function, \(\alpha : \mathcal P \left({X}\right) \rightarrow \mathcal P \left({X}\right)\) such that for all \(S \in \mathcal P \left({X}\right), \alpha (S )= X \backslash S\). Show that \(\alpha\) is a bijection. What is \(( \alpha \circ \alpha) (S)\) for any \(S \in \mathcal P \left({X}\right)?\)

OpenStudy (usukidoll):

WHERE DO I BEGIN BESIDES REALIZING THAT it's a surjection and an injection.

OpenStudy (usukidoll):

so it's a surjection which means that every element of the codomain has at least one preimage

OpenStudy (usukidoll):

and for a surjection for all values of x1 and x2 belonging in X, x1 doesn't equal to x2 and that implies that the f(x1) doesn't equal to f(x2)

OpenStudy (usukidoll):

injection is a one-to-one function

OpenStudy (usukidoll):

all I know is the complement part.. we have x that belongs to X, but not S

OpenStudy (anonymous):

So\[ a(S) = X\setminus S \]

OpenStudy (usukidoll):

Let \(X\) be a set and consider the function, \(\alpha : \mathcal P \left({X}\right) \rightarrow \mathcal P \left({X}\right)\) such that for all \(S \in \mathcal P \left({X}\right), \alpha S = X \backslash S\). Show that \(\alpha\) is a bijection. What is \(( \alpha \circ \alpha) (S)\) for any \(S \in \mathcal P \left({X}\right)?\)

OpenStudy (usukidoll):

I BROKE THE QUESTION FMLLLl

OpenStudy (usukidoll):

yes @wio

OpenStudy (anonymous):

Well \((\alpha \circ \alpha )(S) = X\setminus (X\setminus S)= S\)

OpenStudy (anonymous):

So it is its own inverse.

OpenStudy (usukidoll):

yes but how to prove that... ?!

OpenStudy (usukidoll):

or is it a one liner proof?

OpenStudy (anonymous):

It's really a matter of how much rigor they really want here.

OpenStudy (usukidoll):

let's make it as simple as it can get

OpenStudy (anonymous):

Let's see.... For any S, we want to show there is only one \(X\setminus S\)

OpenStudy (usukidoll):

yeah.. . ... but what a sec by complement def we have an x in X but not in S

OpenStudy (anonymous):

That much is certain. But then we need to show that for any \(X\setminus S\) there is only one corresponding \(S\).

OpenStudy (anonymous):

Since it is so obvious, a proof by contradiction might be easiest.

OpenStudy (anonymous):

If it isn't a bijective function, then that would mean \(X\setminus S\) doesn't exist, or there are two of them.

OpenStudy (usukidoll):

two X's?

OpenStudy (usukidoll):

wait where is the contradiction part.. like we say that the alpha isn't a bijection... so it's not an injection or surjection... no preimage..

OpenStudy (anonymous):

Either not an injection, or not a surjection

OpenStudy (anonymous):

Again, it is hard to say how much rigor is required here.

OpenStudy (anonymous):

Is there a precise definition of injection or surjection we can start off with?

OpenStudy (usukidoll):

yeah there is

OpenStudy (usukidoll):

wait let me latex it and get it back here

OpenStudy (usukidoll):

A function \(f(x) X \rightarrow Y\) with the property \((\forall Y \in Y)( \exists X \in X ) [f(x) = y ]\) is a surjection of \(X\) onto \(Y\).

OpenStudy (anonymous):

you ever use the open study quoting feature?

OpenStudy (usukidoll):

no

OpenStudy (anonymous):

If lets you do this :\(\color{blue}{\text{Originally Posted by}}\) @UsukiDoll A function \(f(x) X \rightarrow Y\) with the property \((\forall Y \in Y)( \exists X \in X ) [f(x) = y ]\) is a surjection of \(X\) onto \(Y\). \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Make multi-latex posts easier to edit.

OpenStudy (usukidoll):

A function \(f: X \rightarrow Y\) with the property \( \forall x_1, x_2 \in X)[ x_1 \neq x_2] \rightarrow f(x_1) \neq f(x_2)\) is an injection of \(X\) into \(Y\)

OpenStudy (usukidoll):

the injection is one to one

OpenStudy (anonymous):

Okay, for surjection, it is easy.

OpenStudy (usukidoll):

alright so the surjection in addition to what I wrote says that every element of the codomain has at least one preimage

OpenStudy (usukidoll):

so the negation would be that there isn't a preimage for every element of the codomain

OpenStudy (anonymous):

In our case, the definition changes to be:\[ \forall T\in \mathcal P(X) \ \exists S\in \mathcal P(X)[X\setminus S = T] \]

OpenStudy (anonymous):

We need to prove this is true.

OpenStudy (usukidoll):

for all T in a power set X there is an S belonging to a power set x ... so x = t

OpenStudy (anonymous):

For all set T that are subsets of X, there is a set S subset of X which is basically the complement.

OpenStudy (usukidoll):

damn if only I could translate these things better... at least I know the symbols... so the complement..we need to prove that

OpenStudy (usukidoll):

gahhhh! err there's a composite function too which is alpha dot alpha OH! that's like the composite of itself like you know g o g means put your g function inside g. so the alpha function is put inside its self and then oh dang I have no idea what's next .

OpenStudy (usukidoll):

and there's power sets too.

OpenStudy (usukidoll):

alright... so our goal is to prove by contradiction that this stupid thing isn't a bijection which means that's it's not a surjection (1 to 1) and an injection (no element has a preimage)... NOW WHERE TO BEGIN ROFL!

OpenStudy (anonymous):

We don't need to actually prove a contradiction. We can do a direct proof since we have this definition.

OpenStudy (usukidoll):

ooo

OpenStudy (dan815):

what is all this mumbo jumbo

OpenStudy (usukidoll):

discrete math @dan815 hop in

OpenStudy (usukidoll):

except this is proving functions.. s(*(*(#)@* . Makes set theory and induction look easy

OpenStudy (anonymous):

First, can we start with: \[ \forall T\in \mathcal P(X)\ \exists S\in \mathcal P(x)\quad X\setminus T = S \]We're just saying there is a subset in \(X\) that is complement for \(T\).

OpenStudy (dan815):

xcool is there an MIT course on it!

OpenStudy (usukidoll):

so for all T belonging in a power set X, there is some S belonging in a power set X such that the complement of X has an S?!

OpenStudy (anonymous):

We are just saying that every T has an complement.

OpenStudy (anonymous):

and that complement is also a subset of the universal set X.

OpenStudy (usukidoll):

alright so it belongs to a universal set X

OpenStudy (anonymous):

Actually, don't even worry about complements. Let's just say you can take the set difference from X and a subset T, and the result is a subset S

OpenStudy (anonymous):

We can start out even more basic.

OpenStudy (anonymous):

We can start with the premise that set minus will yield a result:\[ \forall T\in\mathcal P(X) \quad X\setminus T=S \]We know that \(S\) is in \(\mathcal P(X)\) because it must be a subset of \(X\). This allows us to add \(\exists S\in \mathcal P(X)\).

OpenStudy (anonymous):

Next thing we do is complement both sides\[ \forall T\in\mathcal P(X)\ \exists S\in\mathcal P(S) \quad (X\setminus T)^C=S^C \]Then intersect them both with \(X\)\[ \forall T\in\mathcal P(X)\ \exists S\in\mathcal P(S) \quad X\cap (X\setminus T)^C=X\cap S^C \]

OpenStudy (anonymous):

We can simplify \(X\cap S^C\) to be \(X\setminus S\).

OpenStudy (usukidoll):

huh where did you get the X cap...?! on the left side... I know that for the X\T there's a set property

OpenStudy (anonymous):

It turns out that \[X\cap(X\setminus T)^C = X\cap(X\cap T^C)^C= X\cap(X^C\cup T)= (X\cap X^C)\cup (X\cap T)= T\]

OpenStudy (anonymous):

This gives us \[ \forall T\in\mathcal P(X)\ \exists S\in\mathcal P(S) \quad T=X\setminus S \]Which proves we have a surjection.

OpenStudy (usukidoll):

what about for injection? .-. is it easy to prove or not?

OpenStudy (anonymous):

that one is maybe a candidate for direct or contrapositive.

OpenStudy (usukidoll):

let's try contrapositive

OpenStudy (anonymous):

\[ X\setminus S_1 = X\setminus S_2\implies S_1=S_2 \]

OpenStudy (anonymous):

You would have to prove this.

OpenStudy (anonymous):

It would work the same way. Start with \(X\setminus S_1 = X\setminus S_2\) and then you do \(^C\) and \(X\cap\) to both sides.

OpenStudy (usukidoll):

k I'll finish this when I wake up...super sleepy.

OpenStudy (usukidoll):

did some reading and practices earlier... energy went need sleep

OpenStudy (usukidoll):

uh oh. I got all the way to \[(X \cap X^C) \cup ( X \cap T) = T\] there's a law that's written as \[X \cup \emptyset = X \] that leaves \[X \cup (X \cap T)\] if I use idempotent law \[X \cup X = X \] I'll have \[X \cap T\]... how do I get the T to be single for surjection ?

OpenStudy (usukidoll):

wrong law.. I meant. \[X \cap X^c = \emptyset\]

OpenStudy (usukidoll):

\[\emptyset \cup (X \cap T)\]

OpenStudy (usukidoll):

\[X\cup \emptyset = X\] identity law \[X \cap T \]

OpenStudy (anonymous):

Since \(T\subseteq X\) we know that \(T\cap X = T\)

OpenStudy (anonymous):

Since\(X\) has all the elements that \(T\) has, we know the intersection will not take away any elements from \(T\)

OpenStudy (usukidoll):

Suppose \(\alpha\) is an injection. Definition 5.4.5 states that a function \(f: X \rightarrow Y\) with the property \( \forall x_1, x_2 \in X)[ x_1 \neq x_2] \rightarrow f(x_1) \neq f(x_2)\) is an injection of \(X\) into \(Y\)\\ \(X \setminus S_1 = X\setminus S_2 \rightarrow s_1=s_2\) \\ By complementing both sides, we have:\\ \((X \setminus S_1)^C = (X\setminus S_2)^C \rightarrow (s_1)^C=(s_2)^C\)\\ Then, we intersect both of them with \(X\). \\ \(X \cap (X \setminus S_1)^C = X \cap (X\setminus S_2)^C \rightarrow X \cap (s_1)^C= X \cap(s_2)^C\)\\

OpenStudy (usukidoll):

\[X \cap (X \cap s_1^c)^c\]

OpenStudy (usukidoll):

\[X \cap (X^c \cup S_1)\]

OpenStudy (anonymous):

You can't prove it by supposing it is true.

OpenStudy (usukidoll):

O_)OSDApf

OpenStudy (usukidoll):

\[X \cap X^c \cup X \cap S_1\]

OpenStudy (usukidoll):

\[\emptyset \cup X \cap S_1\]

OpenStudy (usukidoll):

\[X \cap S_1\]

OpenStudy (anonymous):

When you use the contrapositive, you are changing what you have to prove, you are not assuming that it is true.

OpenStudy (anonymous):

Since you a proving a conditional statement, then you get to assume the antecedent is true and then show through a direct proof that the consequent must also be true.

OpenStudy (anonymous):

In this case, we start by assuming that \[ X\setminus S_1 = X\setminus S_2 \]is true, and then show that we arrive at \[ S_1=S_2 \]must be true.

OpenStudy (usukidoll):

oops ._.

OpenStudy (usukidoll):

but that's by complementing and intersecting both sides right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

The point is, you have to prove the conditional relationships, and can't start out on the premise that it is true.

OpenStudy (usukidoll):

so suppose it's not an injection?!

OpenStudy (usukidoll):

then do the complement and intersection things as usual.... but would mean that alpha isn't a bijection after all.

OpenStudy (anonymous):

Make no assumptions.

OpenStudy (anonymous):

Only assume \(X\setminus S_1=X\setminus S_2\)